OpenStudy (tootsi123):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (er.mohd.amir):

since two points are given so it satisfy eq. put values of x and y and find A,B,C get requird eq.

OpenStudy (tootsi123):

So 7x-3y=?

OpenStudy (er.mohd.amir):

7A-3B=C

OpenStudy (er.mohd.amir):

or u can find eq passing throug two points y=mx+c pass it by two points finf m and c

OpenStudy (tootsi123):

Oh okay I got it

OpenStudy (er.mohd.amir):

or y-y1={(y2-y1)/(x2-x1)}(x-x1) u can use it

OpenStudy (tootsi123):

Thank you.